# Hardware test and validation of the angular droop control: Analysis and experiments

### v1

@@ IV-A1 Programmable DC/AC converter system
As shown in Fig. 6, a programmable DC/AC converter system consists of a DC/DC converter, a control unit, a DC/AC converter and an output filter.
  • DC/DC converter: a DC/DC stage, consisting of a boost converter, is deployed to regulate the supplied DC voltage to a desired value. In the test bench, a DC power supply provides a constant DC voltage. The half-bridge power modules are intended for rapid prototyping for power electronics [29]. The adopted control scheme is described in Sec. III-B1.
  • Control unit: the real-time control unit includes an analog front-end to connect the sensor measurements, a dual-core 1 GHz ARM processor to run user-generated code. The control unit is capable of handling fast tasks such as the gate pulse generation for power electronics and enables closed-loop control with a switching frequency up to $250$ kHz [30].
@@
  • DC/AC converter and output filter: the two-level three-phase DC/AC converter system consists of half-bridge modules described in [29]. The output filter consists of an inductance in series with a parasitic resistance. Both are set in parallel with a capacitance [31]. The DC/AC converter system has a rated power of $P_{rated}=15$ kW and a nominal voltage of $230\sqrt{2}\,\mathrm{V}$. Therefore, it is suitable for low voltage applications. The control scheme for the DC/AC converter is described in Sec. III-B2.
The parameter values of the DC/DC and DC/AC converter and grid connection filter are summarized in Table I. The overall system is monitored in real-time by a computer, from which the relevant parameters of the controller as well as the auxiliary systems can be modified during runtime using the respective GUI.
Refer to captioncontrol unitDC/DC converter and DC/AC converterpassive components and output filterDC power supply
Figure 6: Overview of a programmable DC/AC converter system for the hardware experiments at the Energy Lab.
IV-A2 Resistive load
@@
For our laboratory experiment, we consider a light-wall shown in Fig. 7. The light-wall consists of 208 light bulbs. Each has a power consumption of $100\,\mathrm{W}$ at nominal voltage of $230\sqrt{2}\,\mathrm{V}$. The light bulbs are evenly distributed over the three phases and can be individually controlled via a programmable logic controller to reach a user-defined load power. Due to the ability of synchronous switching of the light bulbs, variable load steps can be realized using the automation system of the Energy Lab.
IV-A3 Transmission line replica
The transmission line replicas are deployed to reproduce the relevant physical properties of the transmission system [7]. For our experiments, the resistive and inductive or (RL)-type replicas are used. They consist of a resistance set in series with an inductance. The parameter values of the transmission line of our experimental setup are given in Table II.
Figure 7: Resistive load represented by a light-wall.
IV-B Test scenarios
In the remainder, we consider the following test scenarios for our hardware experiments.
  1. 1.
@@ IV-B Test scenarios
    Scenario I: A single DC/AC converter system in closed-loop with the angular droop control is connected to a resistive load. The goal of this scenario revolves around testing:
@@
    • the black start capabilities, namely the ability to form sinusoidal wave after a major event, e.g., a blackout.
@@
    • the direct and indirect implementation schemes of the angular droop behavior and drawing a comparison between them.
  2. 2.
    Scenario II: Two identical DC/AC converter systems, each in closed-loop with the angular droop are connected to a common resistive load. In this scenario, we test:
    • the frequency synchronization capabilities of the angular droop control.
    • the power sharing capabilities in dependence of the angular droop control gains.
TABLE I: Technical details of the hardware setup
DC/DC converter and its control
Symbol Definition Range Value in S.I.
$V^{nom}_{dc}$ DC voltage >0 750
$V_{b}$ source voltage >0 600
$C_{dc}$ DC capacitance >0 $3\cdot 10^{-3}$
$R_{b}$ parasitic resistance >0 $1\cdot 10^{-3}$
$L_{b}$ boost inductance >0 $2.36\cdot 10^{-3}$
$k_{P}$ P-voltage gain >0 $0.3$
$k_{I}$ I-voltage gain >0 $12$
$k_{BP}$ P-current gain >0 $10$
$k_{BI}$ I-current gain >0 $200$
DC/AC converter and its control
Symbol Definition Range Value in S.I.
$V^{*}$ AC voltage amplitude >0 $230\sqrt{2}$
$C$ AC filter capacitance >0 $1\cdot 10^{-5}$
$L$ AC filter inductance >0 $2.36\cdot 10^{-3}$
$R$ AC filter resistance >0 $1\cdot 10^{-3}$
$G$ load resistance >0 $58.77$
$A$ modulation amplitude >0 $0.8132$
angular droop control
Symbol Definition Range Value in S.I.
$P^{*}$ active power >0 $2880$
$\omega^{*}$ frequency >0 $2\pi 50$
$\alpha$ input effort gain >0 $2000$
$\gamma$ steady state gain >0 $5\cdot 10^{4}$
$k_{VP}$ P-voltage gain >0 $0.05$
$k_{VI}$ I-voltage gain >0 $0.4$
$k_{IP}$ P-current gain >0 $10$
$k_{II}$ I-current gain >0 $240$
V Scenario I: Single converter to load
V-A Scenario description
This scenario consists of a balanced three-phase DC/AC converter connected to the resistive load and supplied by a DC power supply behind a DC/DC converter as shown in Fig. 8. For the DC/DC converter, we follow vector control to provide a constant voltage level for the DC/AC stage. The DC/AC converter is controlled by the angular droop control either directly or indirectly via the switching modulation signal $\overline{u}$. We refer the reader to Section III-B for the details related to the individual control schemes.
@@ V-A Scenario description
V-B Experimental procedure
We start by configuring the lab automation system [7] according to the hardware configuration in Fig. 8. Thereafter, the DC supply together with the DC/DC converter are enabled and the following experiments are started.
  1. 1.
    Black start: From zero voltage conditions, the DC/AC converter is connected to the grid by closing the respective three-phase circuit breaker. The amplitude of the switching modulation signal in (14) is set to its nominal value given by,
    $A=\frac{2\,V^{d}}{V^{nom}_{dc}},$
    where $V^{d}>0$ is the amplitude of the output voltage $v$ in (15) and $V^{nom}_{dc}>0$ is the nominal DC voltage.
  2. 2.
    Step in the load power: Starting from nominal steady state conditions, we increase the load power at $0.2\,\mathrm{s}$. Moreover, we study suitable tuning of the angular droop control by varying the control gains $\alpha>0$ and $\gamma>0$ and studying their effect on the angle and frequency error transients and their induced steady states.
  3. 3.
    Direct vs. indirect implementation: We compare the black start and step in the load experiments for the DC/AC converter controlled by the angular droop control according to the direct with the indirect implementation described in Section III-B2.
V-C Challenge 1: Discretization of the angle dynamics
The angular droop control described in Section III-B needs to conform with the restrictions posed by the hardware prototype, in particular its hardware implementation in discrete-time. For this, we discretize the closed-loop angle dynamics in Eq. (3) using forward Euler method as follows,
$\theta(s+1)$ $=\theta(s)+T_{s}\,u_{d}(s)+\omega^{*}\mathds{1}_{n}$ (18)
$u_{d}(s)$ $=-\frac{1}{2}R^{-1}\,(\Gamma\,(\theta(s)-\theta^{*}(s))+P(\theta(s))-P^{*})$
where $s\in\mathbb{Z}$ is the time step, $T_{s}>0$ is the sampling period and $\omega^{*}>0$ is the nominal frequency. Next, we define the angle error coordinate $\Delta\theta(s)=\theta(s)-\theta^{*}(s)$ where $\theta^{*}\in^{n}$ denotes the nominal angle satisfying,
@@ V-C Challenge 1: Discretization of the angle dynamics
Observe that in (19), the nominal steady state angle vector $\theta^{*}$ grows infinitely. This causes a loss of precision for the stored variable $\theta(s)$ due to the limit of available bytes for single precision. This observation motivates the following solution: Find a mapping of the angles $\theta^{*}$ from n to the $n-$th dimensional torus $\mathbb{T}^{n}$. Note that the sine function appearing in the implementation of the modulation signal $\overline{u}$ both for the direct and indirect schemes, namely in (14) and (17) is $2\pi-$ periodic in the angle $\theta(s)$. We proceed by limiting the values of the nominal angle $\theta^{*}(s)$ as follows,
[…]
@@
$\theta(s+1)=\theta^{*}(s+1)+\Delta\theta(s+1).$
Here, $\Delta\theta(s+1)$ is given by (20). therefore the angles $\theta(s),\,s\in\mathbb{Z}$ remains within feasible numerical bounds.
V-D Experimental Results
@@
  1. 1.
[…]
@@
    Figure 9: Averaged DC voltage $V_{dc}$ in (a), leg inductance current $I_{b}$ in (b), active power $P$ in (c) and AC voltage $v$ in (d) in the black start experiment.
  2. 2.
    Step in the load power:
@@
    Figure 10: Averaged DC voltage $V_{dc}$ in (a), leg inductance current $I_{b}$ in (b), active power $P$ in (c) and AC voltage $v$ in (d) in the load step experiment.
    Figure 11: AC inductance current in the load step experiment
@@
    Following the load step change at $t=0.2$s, the active power overshoots to approximately $1.2\cdot 10^{4}$ and settles at $P^{s}\approx 3800$ at steady state from Fig. 10. This is induced by large rush currents accompanying the change in load power as shown in Figure 11. The DC voltage $V_{dc}$ returns to its nominal value after $0.1\,\mathrm{s}$ due to the integral control action of the inner current loop in (11), whereby the leg inductance current reaches a newly induced steady state $I_{b}^{s}\approx 7$ following the load step. The three-phase voltage amplitude drops slightly during the event and stays around its nominal value at steady state.
    Figure 15 shows a drop in the AC frequency during the load power change, however the frequency returns to its nominal steady state despite the increase in load power. Zero steady state frequency error is a characteristic property of the angular droop control. See also Section II-B.
@@
    In the following, we determine a suitable tuning of the angular droop control by studying the influence of the control gains $\alpha>0$ and $\gamma>0$ on the frequency and angle of the modulation signal following a step in the load at $t=0.2\,$s. As depicted in Fig. 12, an increase in $\alpha$ leads to a higher penalty on the angle and power deviations from their steady state. This translates into a larger rate of change of frequency and smaller nadir. This is in accordance with the observation that, the choice of the gain $\alpha$ affects the transient behavior of the AC frequency $\omega$ and angle $\theta$ in (6). To avoid an overshoot and limit the frequency deviation to an acceptable deviation of $0.8$ Hz [33], we fix the value $\alpha=2000$. Fig. 13 shows that, for decreasing values of $\gamma$, the drop in the angle error at steady state or power-to-angle droop behavior is more significant following (5). This results in an induced steady state angle $\theta^{s}$ given by
@@
[…]
  3. 3.
    Comparison between direct and indirect control: For the control gains $\alpha=2000$ and $\gamma=5\cdot 10^{4}$, Fig. 14 and 15 compare the frequency and angle errors resulting from the direct and indirect implementation for the black start and the load step experiment, respectively. Also in the indirect implementation, the frequency is restored to its nominal value within approximately $0.7\,\mathrm{s}$ and the angles converge to an induced steady state angle $\theta^{s}$ as in Eq. (22). In fact, the major difference between direct and indirect control is in the way the control law (14) relates to the modulation signal representing the main input to the DC/AC converter. In the direct implementation, we assign a sinusoidal wave whose angle is directly determined by the angular droop control. The indirect implementation encodes a nested voltage and current control loops relying on tracking a given AC voltage reference, whose angle is described by the angular droop control. This results in different modulation signals $\bar{u}$, where the control effort solely dependent on the gains $\alpha$ and $\gamma$ in the direct scheme. For the indirect scheme, the control effort is also dependent on the choice of the inner (current) and outer (voltage) control loops, see Table I. For our particular control gain choice, this results in a slower convergence rate to a steady state compared to the direct scheme. This is visible from the duty cycles $d$ in Fig. 16, defined by (10). The duty cycle increases linearly in the direct whereas sub-linearly in the indirect implementation scheme.
    Due to its simpler and more intuitive tuning, we adopt the direct implementation of the angular droop control with $\alpha=2000$ and $\gamma=5\cdot 10^{4}$ in the remainder of our hardware experiments.

### v2

@@ II-C Experimental setup
As shown in Fig. 1, the hardware representation of the programmable DC/AC converter system consists of a DC power supply, DC/DC converter, a control unit, a DC/AC converter and an output filter.
II-C1 Programmable DC/AC converter system
  • DC power supply: emulates the power source of the DC/AC converter system and has a rated power of $15\,\mathrm{kW}$.
  • DC/DC converter: consists of a boost converter which is build using a half-bridge module [12] as well as an inductance of $L_{b}=1.3\,\mathrm{mH}$. The system is controlled to regulate the DC bus voltage at the desired value.
@@ II-C1 Programmable DC/AC converter system
  • DC/AC converter and output filter: the two-level three-phase DC/AC converter system consists of SiC-MOSFET half-bridge modules, see Fig. 3, including an embedded DC-bus of $500\,\mu\mathrm{F}$, as described in [12]. To form the three-phase DC/AC converter, the half-bridge modules are connected on the DC-side to form a common DC-bus, which is supplied by the DC/DC converter with a constant voltage. The output filter consists of an inductance in series with a parasitic resistance for each of the phases. They are set in parallel with a capacitance to form a LC-filter [26]. The DC/AC converter system has a rated power of $P_{rated}=15$ kW and a nominal voltage of $230\sqrt{2}\,\mathrm{V}$. The schematics are shown in Fig. 2
  • Control unit: the real-time control unit [4] executes the user code and controls both the DC/DC and the DC/AC stage during the experiments. The control cycle frequency of $20\,\mathrm{kHz}$ is well above the nominal grid frequency of $50\,\mathrm{Hz}$ such that the control latency has only negligible effect on the presented results.
  • Measurement sensors: To obtain the necessary measurements for the control of the converters and presentation of the results, the following sensors are used:
    • Current sensors [1]: the measurement range of $\pm 50\,\mathrm{A}$ is well suited for this application and the bandwidth of $200\,\mathrm{kHz}$ is large enough to cover the relevant frequency range, considering a switching frequency of $20\,\mathrm{kHz}$ and nominal grid frequency of $50\,\mathrm{Hz}$. Further, the typical sensitivity error of the sensors is $\pm 0.4\,\%$ which is sufficient for the control usecase. The input-referred noise is $0.05\,\mathrm{A}$, which is several orders of magnitude smaller than the operating point during the experiments and therefore has negligible effect on the presented results.
    • Voltage sensors [2]: the measurement range of $\pm 800\,\mathrm{V}$ and the bandwidth of $100\,\mathrm{kHz}$ is sufficient to capture the relevant dynamics, considering the switching frequency of $20\,\mathrm{kHz}$ and nominal grid frequency of $50\,\mathrm{Hz}$. Further, the typical sensitivity error of the sensors is $\pm 0.35\,\%$ which is adequate for the control usecase. The input-referred noise is $1.4\,\mathrm{V}$, which is sufficiently small compared to the operating voltage.
    Since each phase voltages and currents are obtained by individual sensors, the values for the sensitivity and offset differ slightly even after calibration. Such sensor-specific errors can lead to an imbalance in the measured phase currents and voltages, which generates inter-harmonic components in derived values such as the active power.
The parameter values of the programmable DC/AC converter are summarized in Table I.
Figure 2: A three-phase DC/AC converter model under study. Each phase-leg consists of a half-bridge module [12]. The modulation input $\bar{u}$ is obtained either directly (direct control) or via well-known cascaded voltage and current control (indirect control), for control diagrams and further details see Appendix Hardware test and validation of the angular droop control: Analysis and experiments.
Figure 3: Exemplary image of the half-bridge modules [12] contained in the power rack.
II-C2 Resistive load
@@
We consider a light-wall to represent a resistive load, see Fig. 4. Each of the light bulbs has a power consumption of $100\,\mathrm{W}$ at nominal voltage of $230\sqrt{2}\,\mathrm{V}$ and can be individually controlled via a programmable logic controller to set the desired power consumption.
Figure 4: Resistive load represented by incandescent light bulbs.
II-C3 Transmission line replica
The physical transmission line properties are emulated using resistive and inductive type line replicas. Discrete hardware elements are connected in series to provide the desired physical quantities in order to replicate the effect of a given length and type of transmission line [36], see Fig. 5. The parameter values of the transmission line of our experimental setup are given in Table I.
Figure 5: Transmission line replica in our experimental setup.
TABLE I: Technical details of the hardware setup
DC/AC converter and its control
Symbol Definition Value (S.I.)
$V^{*}$ AC voltage amplitude $230\sqrt{2}$
$C$ AC filter capacitance $1\cdot 10^{-5}$
$L$ AC filter inductance $2.36\cdot 10^{-3}$
$R$ AC filter resistance $1\cdot 10^{-3}$
$G$ load resistance $58.77$
$V_{dc}$ DC-link voltage $750$
$f_{sw}$ switching frequency $20\,\mathrm{kHz}$
AA modulation amplitude $0.8132$
Angular droop control
Symbol Definition Value (S.I.)
$P^{*}$ nominal active power $2880$
$\omega^{*}$ nominal angular frequency $2\pi 50$
$\alpha$ input effort gain $2000$
$\gamma$ steady state gain $5\cdot 10^{4}$
$k_{VP}$ P-voltage gain $0.05$
$k_{VI}$ I-voltage gain $0.4$
$k_{IP}$ P-current gain $10$
$k_{II}$ I-current gain $240$
Parameters values of the transmission line
Symbol Definition Value in S.I.
$R_{l}$ line resistance $20\cdot 10^{-3}$
$L_{l}$ line inductance $700\cdot 10^{-6}$
III Scenario I: Single converter to load
Angular droop control operates in a different domain than conventional frequency droop by linking active power to angle rather than frequency. Although angle and frequency are related through integration, this distinction poses challenges for the real-world implementation.
III-A Scenario Goals and Description
@@
To experimentally verify the properties of the angular droop hardware implementation under real-world conditions, we start with Scenario I: a single DC/AC converter system in closed-loop with the angular droop control is connected to a resistive load, as shown in Fig. 6.
With Scenario I we show:
@@
  • black start capabilities, namely the ability to form sinusoidal wave after a major event, e.g., a blackout.
@@
  • influence of direct and indirect implementation schemes on the angular droop behavior and drawing a comparison between them. The modulation input $\bar{u}$ of the DC/AC converter is obtained either directly (direct control) or via well-known cascaded voltage and current control (indirect control), details see in Appendix Hardware test and validation of the angular droop control: Analysis and experiments.
The single-converter test case is crucial for assessing the resilience of the closed-loop converter system against disturbances of varying magnitudes. The ability to withstand small to medium disturbances is particularly relevant in conditions close to nominal operation, while grid-forming capabilities become essential following major disruptions such as blackouts. Comparing direct and indirect implementations shows the compatibility of angular droop with well established control architectures, while preserving its properties.
@@
III-B Challenge: Discretization of the angle dynamics
Analysis
The angular droop control described in Section II, also see Appendix Hardware test and validation of the angular droop control: Analysis and experiments, needs to conform with the restrictions posed by real hardware, in particular its implementation in discrete-time. For this, we discretize the closed-loop angle dynamics in (3) using the forward Euler method as follows,
$\theta(s+1)$ $=\theta(s)+T_{s}\,u_{d}(s)+\omega^{*}\mathds{1}_{n}\,$, (11)
$u_{d}(s)$ $=-\frac{1}{2}R^{-1}\,(\Gamma\,(\theta(s)-\theta^{*}(s))+P(\theta(s))-P^{*})\,$,
where $s\in\mathbb{Z}$ is the time step, $T_{s}>0$ is the sampling period and $\omega^{*}>0$ is the nominal angular frequency. Next, we define the angle error coordinate $\Delta\theta(s)=\theta(s)-\theta^{*}(s)$ where $\theta^{*}\in^{n}$ denotes the nominal angle satisfying
@@ III-B Challenge: Discretization of the angle dynamics
Observe that in (12), the nominal steady state angle vector $\theta^{*}$ grows infinitely. This causes a loss of precision for the stored variable $\theta(s)$ due to the limit of available bytes for single precision.
Proposed solution
The previous observation motivates the following solution. We aim to find a mapping of the angles $\theta^{*}$ from n to the $n$-th dimensional torus $\mathbb{T}^{n}$. Note that the sine function appearing in the implementation of the modulation signal $\overline{u}$ both for the direct (38) and indirect (39) schemes is $2\pi$ periodic in the angle $\theta(s)$. We proceed by limiting the values of the nominal angle $\theta^{*}(s)$ as follows,
$\theta^{*}(s+1)=\theta^{*}(s)+T_{s}\,\omega^{*}(\mathrm{mod\,2\pi}),$ (14)
@@
$\theta(s+1)=\theta^{*}(s+1)+\Delta\theta(s+1).$ (15)
Here, $\Delta\theta(s+1)$ is given by (13). Therefore the angles $\theta(s),\,s\in\mathbb{Z}$ remains within feasible numerical bounds. Our proposed solution optimizes the space complexity of hardware implementation while ensuring ease of understanding and practicality.
III-C Experimental Results
@@
III-C1 Black start capabilities
From Fig. 7, we observe that the DC capacitor voltage $V_{dc}$ reaches its nominal value $V^{nom}_{dc}=750$ within $0.4\,\mathrm{s}$ following initial transients. This corresponds to the leg inductance current converging to the steady state value $I^{s}_{b}\approx 5$. Following the transients, the visible voltage ripple is due to the sizing of the DC-link capacitors and remains within an acceptable range for the given resistive load. On the AC side, the active power converges to its nominal value $P^{*}=2880$ which corresponds to a sinusoidal balanced three-phase signal of the output capacitor voltage $v$ with nominal amplitude $V^{d}=230\sqrt{2}$. In essence, Figure 12 shows that both the angle and frequency error converge to zero, i.e., the AC frequency and the phase angle of the modulation signal are at their nominal values, namely $\omega^{*}=2\pi\,50$ and $\theta^{*}(t)=2\pi\,50\,\mathrm{t}$, respectively. Thus, our results demonstrate that, even if we start from initial operating conditions far away from nominal operation, resulting from large disturbances, e.g., a black start, the angular droop control is able to form sine waves rotating at a nominal frequency with desired angle and amplitude and is therefore grid-forming.
@@ III-C1 Black start capabilities
Figure 7: Averaged DC voltage $V_{dc}$ in (a), DC/DC leg inductance current $I_{b}$ in (b), active power $P$ in (c) and AC voltage $v$ in (d) in the black start experiment.
III-C2 Step in the load power
@@
Figure 8: Averaged DC voltage $V_{dc}$ in (a), DC/DC leg inductance current $I_{b}$ in (b), active power $P$ in (c) and AC voltage $v$ in (d) in the load step experiment.
@@
Proceeding from the steady state after the black start, we study a load step change at $t=0.2\,\mathrm{s}$. The active power overshoots to approximately $1.2\cdot 10^{4}$ and settles to a new steady state $P^{s}\approx 3800$, see Fig. 8. The overshoot is caused by large rush currents accompanying the change in load power. The DC voltage $V_{dc}$ returns to its nominal value after $0.1\,\mathrm{s}$ due to the integral control action of the DC voltage controller, whereby the DC/DC leg inductance current reaches a newly induced steady state $I_{b}^{s}\approx 7$ following the load step. The three-phase voltage amplitude drops slightly during the event and recovers to its nominal value at steady state. Figure 13 shows a drop in the AC frequency during the load power change, however the frequency returns to its nominal steady state despite the increase in load power. This zero steady state frequency error is an advantageous, intrinsic property of angular droop control as previously discussed in Sec. II-B and makes the deployment of secondary control to restore the frequency to its nominal value obsolete. Whereas, in a frequency droop controlled setup such a step in the load power causes permanent frequency deviation, unless the power reference is adapted by the secondary control layer. This is shown in Fig. 9, where angular droop is compared to frequency droop upon a load step. Note that the steady state deviation for the frequency droop depends on the chosen droop constant.
Figure 9: Comparison of angular droop and frequency droop for the load step experiment. Angular droop control uses $\alpha=2000,\,\gamma=5\cdot 10^{4}$ and frequency control a $5\%$ droop constant.
@@
In the following, we determine a suitable tuning of the angular droop control by studying the influence of the control gains $\alpha,\gamma>0$ on the frequency and angle of the modulation signal following a step in the load at t=0.2t=0.2\,s. As depicted in Fig. 10, an increase in $\alpha$ leads to a higher penalty on the angle and power deviations from their steady state. This translates into a larger rate of change of frequency and smaller nadir. This is in accordance with the observation that, the choice of the gain $\alpha$ affects the transient behavior of the AC frequency $\omega$ and angle $\theta$ in (6). To avoid an overshoot and limit the frequency deviation to an acceptable deviation of e.g., $0.8$ Hz [7], we fix the value $\alpha=2000$. Figure 11 shows that, for decreasing values of $\gamma$ the power-to-angle droop behavior is more pronounced resulting in larger steady state angle deviations, see (4). The induced steady state angle $\theta^{s}$ is given by
@@
Figure 10: Frequency (a) and angle errors (b) with $\gamma=5\cdot 10^{4}$ and different values for $\alpha\in~\{500,\,1000,\,2000\}$ in the load step experiment.
Figure 11: Frequency (a) and angle errors (b) with $\alpha=2000$ and different values for $\gamma\in\{5\cdot 10^{4},\,5\cdot 10^{5},\,5\cdot 10^{6}\}$ in the load step experiment.
III-C3 Comparison between direct and indirect control
For the hardware implementation, the angle dynamics resulting from the angular droop control can be translated to a modulation input $\bar{u}$ either directly (direct control) or via well-known cascaded voltage and current control (indirect control). For more details see Appendix Hardware test and validation of the angular droop control: Analysis and experiments. For the control gains $\alpha=2000$ and $\gamma=5\cdot 10^{4}$, Fig. 12 and 13 compare the frequency and angle errors resulting from the direct and indirect implementation for the black start and the load step experiment, respectively. In Table II numerical values for the comparison are presented. Using the indirect implementation, the frequency is restored to its nominal value within approximately $0.61\,\mathrm{s}$ and the angles converge to an induced steady state angle $\theta^{s}$ as in Eq. (16). In fact, the major difference between direct and indirect control is in the way the control law (38) relates to the modulation signal representing the main input to the DC/AC converter. In the direct implementation, we assign a sinusoidal wave whose angle is directly determined by the angular droop control. The indirect implementation consists of cascaded voltage and current control loops relying on tracking a given AC voltage reference, whose angle is described by the angular droop control. This results in different modulation signals $\bar{u}$, where the control effort solely dependent on the gains $\alpha$ and $\gamma$ in the direct scheme. For the indirect scheme, the control effort is also dependent on the choice of the current and voltage control loops, see Table I. This is visible from the duty cycles $d_{AC}$ in Fig. 14, defined by (37). The duty cycle increases linearly in the direct whereas sub-linearly in the indirect implementation scheme. For our particular control gain choice, this results in a slower convergence rate to a steady state compared to the direct scheme for the load step experiment, see Table II. Due to its simpler and more intuitive tuning, we adopt the direct implementation of the angular droop control with $\alpha=2000$ and $\gamma=5\cdot 10^{4}$ in the remainder of our hardware experiments.